Suppose the people living in a city have a mean score of

41

and a standard deviation of

3

on a measure of concern about the environment. Assume that these concern scores are normally distributed. Using the

​50%minus−​34%minus−​14%

​figures, approximately what percentage of people have a score​ (a) above

41​,

​(b) above

44​,

​(c) above

35​,

​(d) above

38​,

​(e) below

41​,

​(f) below

44​,

​(g) below

35​,

and​ (h) below

38?
Answer #1

Answer)

When the scores are normally distributed

Then,

According to the emperical rule

If the data is normally distributed

Then 68% lies in between mean - s.d and mean + s.d

95% lies in between mean - 2*s.d and mean + 2*s.d

99.7% lies in between mean - 3*s.d and mean + 3*s.d

Given mean = 41

S.d = 3

Note: normal distribution is symmetrical in nature that is equal data lies on both sides of mean

A)

above 41 = 50%

B)

Above 44

44 is one s.d above the mean

That is in between 38 and 44

68% lies

So, above 44 = (100-68)/2 = 16% lies

C)

above 35

35 = mean - 2*s.d

47 = mean + 2*s.d

So 95% lies in between 35 and 47

So, above 35 = 95 + (5/2) = 97.5% lies

D)

above 38 = 68 + (32/2) = 84%

E)

Below 41 = 50%

F)

Below 44 = 68 + 16 = 84% (same logic used in part D)

G)

Below 35 = 2.5% {as above 35 = 97.5% so below 35 = 100 - 97.5)

H)

Below 38 = 16% { above 38 = 84%, so below 38 = 100-84)